Each of the equal sides of an isoceles (two sides equal) triangle is 2 times the third side. The perimeter of the triangle is 20m. Find the sides of the triangle.
Answer by checkley77(12569)

2X+2X+X=20
5X=20
X=20/5
X=4 LENGTH OF THE THIRD SIDE
2*4=8 LENGTH OF THE 2 EQUAL SIDES.
PROOF:
8+8+4=20
20=20